OpenStudy (pooh):

If you graph (0,56) and (56,0) where will they intercept.

OpenStudy (anonymous):

since those are points, they probably wont.

OpenStudy (anonymous):

but if you are talking about lines, then they will intersect at (56,56) think about it graphically. if one is a straight line at 56x and the other a straight line at 56y, then every point along those lines will have the same x or y value. they will meet when those values are the same

OpenStudy (anonymous):

56-0 ----- 0-56 56/-56 =-1 =slope y-0 ---- x-56 y=(x-56)-1 y=-x+56 points are the same

OpenStudy (pooh):

so math frick are you saying that the answer is -1

OpenStudy (anonymous):

(0,56) is a point up the y-axis and the other one is a point along the x-axis. What do you want to do with the points?

OpenStudy (anonymous):

the slope is -1

OpenStudy (pooh):

The question asked me with those numbers where would the line intersect

OpenStudy (anonymous):

I you join up the points to make a line , they intersect the axes at those same points.

OpenStudy (anonymous):

you should try sketching it on a graph paper for better understanding